D= 7i + 9j - 4k




E= 8i + 9j + 5k





Calculate D × E




Find the angle between d and e




Show that D × E is perpendicular to both D and E

1
Expert's answer
2022-03-02T10:24:52-0500
"{\bf D}\times {\bf E}=( 7i + 9j - 4k)\times (8i + 9j + 5k)\\
=63k-35j-72k+35i-32j+36i\\
=71i-67j-9k"

"\theta=\sin^{-1}\frac{\sqrt{71^2+67^2+9^2}}{\sqrt{6^2+9^2+4^2}\sqrt{8^2+9^2+5^2}}"

"\theta=41^\circ"